OpenStudy (anonymous):

Which equation represents the following word problem “When a number n is divided by 3 then increased by 4, the result is 7”? 3n – 4 = 7 n over three - 4 = 7 n over three + 4 = 7 3n + 4 = 7

ganeshie8 (ganeshie8):

we simply do wat the question says

ganeshie8 (ganeshie8):

number n is divided by 3 :- \(\large \frac{n}{3}\)

ganeshie8 (ganeshie8):

then increased by 4 :- \(\large \frac{n}{3} + 4\)

ganeshie8 (ganeshie8):

the result is 7 :- \(\large \frac{n}{3} + 4 = 7\)

ganeshie8 (ganeshie8):

see if that makes some sense